# Surface Tension of a Topological Phase

### v2

@@ VI Role of Slow Quench
In the main text we had shown the results for a sudden quench. We now analyze the growth and shrinkage of clusters under slow quench protocol of the external magnetic field $h(t)$ where we consider
$h(t)=\begin{cases}-h_{f}+\frac{2h_{f}t}{\tau},\text{~{}~{}~{}~{}for $0<t<\tau$,}\\ ~{}~{}~{}h_{f},\text{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}for $t\geq\tau$.}\end{cases}$ (S33)
@@
Thus, $\tau$ is the timescale of the quench (see Fig. S9(a)) over which $h(t)$ changes from $-h_{f}$ to $h_{f}$ (where we assume $h_{f}>0$) and the limit $\tau=0$ corresponds to the sudden quench considered in the main text. If $t$ and $\tau$ are chosen in the units of Monte Carlo steps per site, then we find that the size of a cluster having up-spins at a fixed time $t>\tau$ decreases with the increase of $\tau$ for both classical ($\kappa=0$) and topological ($20<|\kappa|<$ 2) situations, as can be seen from average local magnetization $\langle s_{x,y}\rangle$ in Fig. S9(b,c). This observation can be explained as follows: in this slow quench protocol, when $t<\tau/2$, $h(t)<0$ and thus a cluster having up-spins is not in the global free-energy minima and reduction of its size (i.e. growth of its larger background region) is more favored. Later, when $h(t)>0$ (for $t>\tau/2$), the cluster having up-spins reaches the global free-energy minima and starts showing growth (shrinkage) determined by usual nucleation physics. However, due to the reduction of its size when $t<\tau/2$, the size of a cluster at a fixed time $t>\tau$ becomes smaller compared to that for sudden quench situation ($\tau=0$). Thus, if a cluster grows under a sudden quench, it can be made to shrink when the quench protocol is changed to a slow quench performed over a time $\tau$, since the nucleation phenomenon is decided by the adiabatic limit where the instantaneous free-energy landscape determines the nucleation process. This leads to an increase in the effective critical cluster size with increasing $\tau$ in the slow quench protocol. Thus, the timescale $\tau$ can effectively tune the critical cluster size and influence the growth (shrinkage) of clusters.

### v3

@@ Appendix E Role of Slow Quench
We now analyze the growth and shrinkage of clusters under slow quench protocol of the external magnetic field $h(t)$ where we consider
$h(t)=\begin{cases}-h_{f}+\frac{2h_{f}t}{\tau},\text{~~~~for $0<t<\tau$,}\\ ~~~h_{f},\text{~~~~~~~~~~~~~for $t\geq\tau$.}\end{cases}$ (43)
@@
Thus, $\tau$ is the timescale of the quench (see Fig. 14(a)) over which $h(t)$ changes from $h_{f}$ to $h_{f}$ (where we assume $h_{f}>0$) and the limit $\tau=0$ corresponds to the sudden quench considered in the main text. If $t$ and $\tau$ are chosen in the units of Monte Carlo steps per site, then we find that the size of a cluster having up-spins at a fixed time $t>\tau$ decreases with the increase of $\tau$ for both classical ($\kappa=0$) and topological ($0<|\kappa|<2$) situations, as can be seen from average local magnetization $\langle s_{x,y}\rangle$ in Fig. 14(b,c). This observation can be explained as follows: in this slow quench protocol, when $t<\tau/2$, $h(t)<0$ and thus a cluster having up-spins is not in the global free-energy minima and reduction of its size (i.e. growth of its larger background region) is more favored. Later, when $h(t)>0$ (for $t>\tau/2$), the cluster having up-spins reaches the global free-energy minima and starts showing growth (shrinkage) determined by usual nucleation physics. However, due to the reduction of its size when $t<\tau/2$, the size of a cluster at a fixed time $t>\tau$ becomes smaller compared to that for sudden quench situation ($\tau=0$). Thus, if a cluster grows under a sudden quench, it can be made to shrink when the quench protocol is changed to a slow quench performed over a time $\tau$, since the nucleation phenomenon is decided by the adiabatic limit where the instantaneous free-energy landscape determines the nucleation process. This leads to an increase in the effective critical cluster size with increasing $\tau$ in the slow quench protocol. Thus, the timescale $\tau$ can effectively tune the critical cluster size and influence the growth (shrinkage) of clusters.